OpenStudy (anonymous):

Two legs of a right triangle are equal and have length a, which is an integer. The hypotenuse of that triangle has length c, which is also an integer.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

is it always sometimes or never true

OpenStudy (anonymous):

Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

ive heard of it

OpenStudy (anonymous):

It says that if a and b are the lengths of two legs of a right triangle and c is the length of the hypotenuse, then \[a ^{2} + b ^{2} = c ^{2}\]

OpenStudy (anonymous):

Now if a and be are equal, this becomes \[a ^{2} + a ^{2} = b ^{2}\] or 2a^2 = b^2.

OpenStudy (anonymous):

Am I making sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So... if we take the square root of both sides, we get \[a \sqrt{2} = b\] Now the square root of 2 is definitely not an integer, and if you multiply the square root of 2 by an integer, you won't get another integer. So therefore, the answer is it's never true. Does that help?

OpenStudy (anonymous):

yes thank you very much